the ? is determine whether is it posssible to draw a triangle with sides of the given measures.write yes or no and the promble is 15,12,9
Answer by stanbon(75887)

determine whether is it posssible to draw a triangle with sides of the given measures.write yes or no and the promble is 15,12,9.
If it is a right triangle 15^2 = 12^2+9^2
225=144+81
225=225
So, it is a right triangle.
Cheers,
Stan H.